OpenStudy (anonymous):

What is an equation for a circle if the endpoints of a diameter are at (3, 4) and (-3, -2)? Then graph the circle.

OpenStudy (freckles):

find the midpoint

OpenStudy (freckles):

that will be your center

OpenStudy (freckles):

then find the distance between those points that will be length of the diameter than half it to get the length of the radius

OpenStudy (anonymous):

midpoint is (0,1)?

OpenStudy (freckles):

yep

OpenStudy (freckles):

\[(x-h)^2+(y-k)^2=r^2\] so you found (h,k)=(0,1)

OpenStudy (freckles):

you could use the distance formula as I said earlier or you can use this formula and plug in either point and solve for r

OpenStudy (triciaal):